OpenStudy (anonymous):

Which is a counterexample to this conjecture? The sum of 3 and any other number is an odd number. A. 3 + 4 = 7 B. 3 + 51 = 54 C. 3 + 58 = 61 D. 3 + 98 = 101

OpenStudy (anonymous):

I'm thinking it's either between A or B. @No.name

OpenStudy (anonymous):

i don't know what is counter-example :(

OpenStudy (anonymous):

OH okay! That's fine.

OpenStudy (anonymous):

@tester97 @Whitemonsterbunny17

OpenStudy (zzr0ck3r):

Its not A. The claim says that if we add any number to 3 we get an odd number. Is this true? How can we show that it is not? In what example to they add a number to 3, and do NOT get an odd number. (i.e. they get an even number)?

OpenStudy (anonymous):

Is it C

OpenStudy (anonymous):

Oh so basically , a counter-example is one that rules out the statement

OpenStudy (zzr0ck3r):

is 3+58 =61 an even number?

OpenStudy (zzr0ck3r):

we are trying to show that it is NOT ALWAYS the case that when we add a number to 3 we get an even number.

OpenStudy (anonymous):

oh i mean D

OpenStudy (anonymous):

wait thats an odd numbr

OpenStudy (zzr0ck3r):

do you know what even and odd is @VivaLaChloe1 ?

OpenStudy (anonymous):

No duhh. .-.

OpenStudy (zzr0ck3r):

Ok then. Again, in which example do they add a number to 3 and get an EVEN number?

OpenStudy (anonymous):

So is it B

OpenStudy (zzr0ck3r):

correct. Do you see how that is the example that proves that if we add a number to 3, we dont always get an odd number?

OpenStudy (anonymous):

Kinda

OpenStudy (zzr0ck3r):

we added a number to 3 and got an even number right?

OpenStudy (zzr0ck3r):

thus it is NOT the case the 3 plus a number is always odd.

OpenStudy (zzr0ck3r):

any of these would have been a counter example 3+3=6 3+5=8 3+11=14 3+any odd number = even number

OpenStudy (zzr0ck3r):

yes @No.name it is just an example to show something is not true if we have the claim that all animals are dogs then a counter example would be a cat. because a cat IS an animal that IS NOT a dog.

OpenStudy (anonymous):

oh i see,